OpenStudy (anonymous):

Identify the 12th term of a geometric sequence where a1 = 8 and a6 = -8,192. PLS HELP. Will fan and give medal. :)

OpenStudy (anonymous):

A. 134,217,728 B. 33,554,432 C. -33,554,432 D. -134,217,728

OpenStudy (campbell_st):

well before you find the term, you need the common ratio r a term in a geometric series is \[a_{n} = a_{1} \times r^{n - 1}\] so use the 6th term, n = 6 and the 1st term to find r \[-8192= 8 \times r^{6 - 1}~~~~~or~~~~~~~ -8192 = 8 \times r^5\] solve for r when you get it substitute back into the 1st formula with n = 12 and a1 = 8 to find the 12th term hope it helps

OpenStudy (anonymous):

I got C for the answer.

OpenStudy (campbell_st):

C was my answer... well done

OpenStudy (anonymous):

Great, thank you so much for all your help!